OpenStudy (anonymous):

Method: Exponential Model Q=Qoe^kt How long does it take $1125 to triple if it is invested at 7% interest, compounded quarterly?

OpenStudy (anonymous):

3 years

OpenStudy (anonymous):

j

OpenStudy (anonymous):

wrong.. the answer is 15.8 years... but when I do it I get 15.6y.. dont know why...

OpenStudy (nincompoop):

what is o?

OpenStudy (anonymous):

Qo is initial quantity... in this case 1125

OpenStudy (anonymous):

did you add the 2 at the end like you do when it does more than dobel

OpenStudy (nincompoop):

what is your k?

OpenStudy (anonymous):

No, what i dont know how to do is the "quarterly compunded thing" i dont know how to place it in the equation...

OpenStudy (anonymous):

@nincompoop k=0.07

OpenStudy (nincompoop):

what is your k?

OpenStudy (nincompoop):

do you know how to solve for t?

OpenStudy (anonymous):

kinda... this is my work: 3(1125) = 1125e^0.07t ln3 = lne^0.07t 1.0986 = 0.07t 15.67 = t But the correct answer is supposed to be 15.8years... there is something wrong that I cant see... and it has to do with the quarterly compound part, but I dont know how to palce that on the equation.

OpenStudy (nincompoop):

where did you get 3 from?

OpenStudy (anonymous):

3(1125) = 1125e^0.07t 3375 = 1125e^0.07t divide by 1125 both side... 3 = e^0.07t

OpenStudy (anonymous):

@Hero @ganeshie8 @wio @Luigi0210 @beccaboo333 @eliassaab does any of u guys know exponential model problems? pls I need to understand this for my math test tomorrow :)

OpenStudy (nincompoop):

there might be a problem with the equation

OpenStudy (anonymous):

no the equation is right... I know where the problem is at.. I jsut dont know how to place it in the equation.... the problem says that the 7% compounds quaterly, (4) ... I jsut dont know how to place that 4 because everywhere I do it just gets wrong... idk if I have to use the Half life formula or what.. :(

OpenStudy (anonymous):

@nincompoop

OpenStudy (nincompoop):

hehe I deleted it, ganeshie

OpenStudy (nincompoop):

I told him that the equation is wrong, but he said it's correct so

ganeshie8 (ganeshie8):

e = continuous but QUARTERLY is ur compounding - so formula changes

OpenStudy (anonymous):

mmm.. @nincompoop im sorry for not trusting u i just cant make sense of it i still dont get it tho :/

OpenStudy (anonymous):

why 1+?

OpenStudy (nincompoop):

there are two compounding formulas one is continuous, which makes use of e and the other is by terms like annually, biannually, quarterly, monthly, which is the one ganeshie used

OpenStudy (irishboy123):

Use t = 4*3 = 12 (quarters as compounds quarterly) and 7/4% (quarterly rate) and see where that gets you!! Then try the rate as r = [(1+7%)*(1/4)] - 1 to further refine if you are still not bang on.

OpenStudy (nincompoop):

it requires a proof why 1+ ... but all you need to know for now is when to use one over the other

OpenStudy (anonymous):

so what would be the equation for it? like the formula.... Q=Qo1+K^t ? @nincompoop

OpenStudy (anonymous):

@nincompoop done thx u soo much :) the best